# Discovering a Novel Dynamics Mechanism for Charmonium Scattering

### v1

@@ Abstract
In this letter, we propose a novel mechanism for charmonium scattering that utilizes the internal structure of charmonium. By capturing light flavor quarks and anti-quarks from the vacuum, charm and anti-charm quarks form virtual charmed mesons, which mediate an effective one-boson exchange process. This approach accurately reproduces the di-$J/\psi$ invariant mass spectrum observed by CMS and LHCb, demonstrating its validity. Our mechanism offers a comprehensive framework for understanding charmonium scattering and is applicable to the scattering problems involving all fully heavy hadrons, an area of increasing interest.
@@
Introduction.—Scattering problems, as a fundamental topic in physics, offering critical insights into the interactions and internal structures of matter. Using this approach, Yukawa predicted the pion meson Yukawa:1935xg , which mediates the nuclear force, establishing a cornerstone of modern nuclear physics. In this century, advances in experimental techniques have led to the discovery of numerous new hadronic states, presenting unique opportunities to explore hadron scattering. Noteworthy achievements include the identification of exotic molecular states such as $P_{c}$ states LHCb:2015yax ; LHCb:2019kea and $T_{cc}(3875)^{+}$  LHCb:2021vvq ; LHCb:2021auc . These discoveries not only mark the advent of ’Particle 2.0’ but also enhance our understanding of the nonperturbative aspects of the strong force Chen:2016qju ; Richard:2016eis ; Chen:2016spr ; Lebed:2016hpi ; Esposito:2016noz ; Guo:2017jvc ; Ali:2017jda ; Brambilla:2019esw ; Liu:2019zoy ; Chen:2022asf .
@@
Recently, the CMS Collaboration reported a high-precision measurement of the invariant mass spectrum of di-$J/\psi$ CMS:2023owd , confirming the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealing additional enhancements in the di-$J/\psi$ invariant mass spectrum. Furthermore, ATLAS has joined this effort, alongside LHCb and CMS ATLAS:2023bft . This novel finding surprised the community, prompting various explanations, including exotic hadronic states liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv and coupled-channel dynamics generation Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . However, a substantial issue—charmonium scattering—has not been fundamentally resolved, despite some tentative efforts Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . The highly effective one-boson ($\pi$, $\sigma$, $\rho$, $\cdots$) exchange mechanism for describing hadron scattering fails in quantitatively characterizing charmonium scattering, as charmonium is a typical example of a fully heavy hadron. Thus, studying how charmonium undergoes scattering has become a particularly typical issue in current scattering problems and urgently requires a solution.
@@
Figure 1: A novel approach enabling one-boson exchange for effective charmonium scattering. The cross symbol on the left indicates that direct light flavor one-boson exchange ($\pi,\rho,\sigma,~{}\cdots$) cannot occur.
@@
In this letter, we propose a novel mechanism for charmonium scattering. As illustrated in Fig. 1, the charm and anti-charm quarks within charmonium can capture light flavor anti-quark and quark, respectively, created from the vacuum to form virtual charmed mesons. These virtual mesons provide a source that enables one-boson exchange to be effective for charmonium scattering. A direct application of this novel mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb, demonstrating the mechanism’s validity. We should emphasize that this mechanism is not limited to charmonium scattering; it can also be applied to investigate the scattering problems involving all fully heavy hadrons, a topic that is gaining attention from the community, including lattice research groups Lyu:2021qsh ; Mathur:2022ovu .
@@
Application to reproduce the CMS and LHCb data.—The di-$J/\psi$ invariant mass spectrum observed by the CMS and LHCb Collaborations exhibit significant features around the dip near 6.8 GeV and the peak near 6.9 GeV, corresponding closely to the $\psi\psi(3686)$ and $\psi\psi(3770)$ thresholds, respectively. In contrast, the bumps near 6.6 GeV and 7.3 GeV do not align with any thresholds involving particles from the $J/\psi$ family. Therefore, to capture the main features, we consider couplings among the three channels: ${J/\psi J/\psi,J/\psi\psi(3686),J/\psi\psi(3770)}$.
@@
When one-boson exchange is introduced at the quark level, as depicted in Fig. 1, the scattering mechanism between two charmonia can proceed as shown in Fig. 2. Initially, the two charmonia, $\psi_{1}$ and $\psi_{2}$, couple into pairs of virtual $D$ mesons. Within each pair, one $D$ meson acts as a spectator, while the other engages in an interaction mediated by the exchange of a light hadron, $\mathcal{P}$. Subsequently, the interacting virtual $D$ mesons combine with the spectators to form the final charmonia, $\psi_{3}$ and $\psi_{4}$.
@@
As shown in Fig. 2, we have neglected the contributions from charmed-strange meson loops. This is primarily due to the significantly lower creation rate of $s\bar{s}$ quark pairs compared to $u\bar{u}$ or $d\bar{d}$ quark pairs, with a ratio of approximately ${1}/{\sqrt{3}}$. Consequently, contributions from charmed-strange mesons are at least ${2}/{3}$ smaller. Additionally, the higher threshold for producing charmed-strange meson pairs leads to further suppression by the loop function. Therefore, to focus on the primary physics, we exclude contributions from charmed-strange meson loops.
Furthermore, Fig. 2 illustrates that the exchanged light hadron, $\mathcal{P}$, is limited to $\sigma$ and $\eta$. This restriction arises because, considering the loop mechanism as a black box, the interaction between two charmonia and a light hadron must comply with the conservation laws of the strong interaction. Given that the mass of the exchanged particle should be as small as possible, only $\sigma$ and $\eta$ are viable candidates.
We perform the entire calculation by adopting the effective Lagrangian approach, with all relevant Lagrangians being Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Burdman:1992gh ; Casalbuoni:1996pg
@@
Here, $v\equiv i\overleftrightarrow{\partial}/2\sqrt{m_{i}m_{f}}$, with $m_{i,f}$ representing the masses of the initial or final heavy meson Ding:2021igr . For the coupling constants $g$, $f_{\pi}$, and $g_{s}$, we follow Ref. Ding:2021igr ; Falk:1992cx ; Isola:2003fh ; Liu:2009qhy ; Chen:2019asm and set them as $g=0.59$, $f_{\pi}=132$ MeV, and $g_{s}=0.76$, respectively. The coupling constants $g_{D^{\ast}D^{\ast}\psi}$, $g_{DD\psi}$, and $g_{D^{\ast}D\psi}$, according to heavy quark effective theory, can be related to the parameter $f_{\psi}$ as $\frac{g_{D^{\ast}D^{\ast}\psi}}{m_{D^{\ast}}}=\frac{g_{DD\psi}}{m_{D}}=g_{D^{\ast}D\psi}=\frac{m_{\psi}}{m_{D}f_{\psi}}$.
@@
×$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
×$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
@@
×$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
×$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
@@
where $\mathcal{V}_{\gamma\delta}$ denotes the Lorentz structure of the loop integral, $\mathcal{F}_{q}(q)$ is the form factor for the exchanged light hadron $\mathcal{P}$ to account for the off-shell effect, and $\mathcal{F}_{q_{i}}(q_{i})$ is a form factor used to avoid ultraviolet divergence in the loop integral. We use the widely adopted monopole form for $\mathcal{F}_{q}(q)$, given by $\mathcal{F}_{q}(q)=\frac{\Lambda_{q}^{2}-m_{q}^{2}}{\Lambda_{q}^{2}-q^{2}}$, where $\Lambda_{q}$ is parametrized by a free parameter $\alpha_{q}$ as $\Lambda_{q}=m_{q}+\alpha_{q}\Lambda{\mathrm{QCD}}$, with $\Lambda{\mathrm{QCD}}=220$ MeV and $\alpha_{q}$ is an adjustable parameter. For $\mathcal{F}_{q_{i}}(q_{i})$, to avoid possible unphysical poles on the real axis of the momentum sheet, we adopt $\mathcal{F}_{q_{i}}(q_{i})=\exp(-\vec{q}_{i}^{~{}2}/\Lambda_{i}^{2})$.
Additionally, the reason for multiplying an overall factor of 16 in both Eq. (6) and Eq. (7) is that each $D$ meson loop in Fig. 2 can involve not only $D^{(\ast)}D^{(\ast)}\bar{D}^{(\ast)}$ configurations but also $\bar{D}^{(\ast)}\bar{D}^{(\ast)}D^{(\ast)}$ configurations. Furthermore, for both of these cases, there are charged and neutral configurations. Thus, there are a total of 4 configurations for each loop, resulting in 16 configurations for each diagram in Fig. 2, all of which have the same value.
Using Eqs. (6)-(7), we can express the effective potential in momentum space as
$\mathcal{V}_{(i)}(\sqrt{s},\vec{k},\vec{k}^{\prime})=-\frac{\mathcal{M}_{(i)}(\sqrt{s},\vec{k},\vec{k}^{\prime})}{2\sqrt{p_{1}^{0}p_{2}^{0}p_{3}^{0}p_{4}^{0}}}.$ (8)
Here, $(i)$ stands for $(a)$ or $(b)$ given in Fig. 2, $\sqrt{s}$ is the center-of-mass energy, $\vec{k}$ and $\vec{k}^{\prime}$ are the three-momenta of the initial and final particles in the center-of-mass frame as $\vec{k}=(0,0,k)$ and $\vec{k}^{\prime}=(k^{\prime}\sin\theta,0,k^{\prime}\cos\theta)$, respectively, and $p_{j}^{0}$ denotes the energy of particle $j$.
@@
In this letter, we consider only the $S$-wave behavior in the di-$J/\psi$ spectrum. Therefore, we first perform a partial wave expansion, which can be calculated as
$\mathcal{V}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})\sim 2\pi\int d\cos\theta d_{\lambda\lambda^{\prime}}^{J}(\theta)\mathcal{V}_{\lambda,\lambda^{\prime}}(\sqrt{s},\vec{k},\vec{k}^{\prime})$ (9)
@@
with $k^{(\prime)}=|\vec{k}^{~{}(\prime)}|$, where $\lambda$ and $\lambda^{\prime}$ is the helicity indices of the initial and final states, respectively, and $d_{\lambda,\lambda^{\prime}}^{J}(\theta)$ is the Wigner-$d$ matrix.
@@
Next, we substitute the obtained partial wave effective potentials into the coupled-channel Lippmann-Schwinger equation
$T^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})$ $=$ $\mathcal{V}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})+\sum_{\lambda^{\prime\prime}}\int\frac{k^{\prime\prime 2}dk^{\prime\prime}}{(2\pi)^{3}}$ (10)
×$\times\mathcal{V}^{J}_{\lambda,\lambda^{\prime\prime}}(\sqrt{s},k,k^{\prime\prime})G_{0}(\sqrt{s},k^{\prime\prime})$
×$\times T^{J}_{\lambda^{\prime\prime},\lambda^{\prime}}(\sqrt{s},k^{\prime\prime},k^{\prime}),$
@@
then, we can extract the on-shell coupled-channel $T$-matrices $T^{J}(\sqrt{s})$. Following a similar approach as in Refs. Gong:2020bmg ; Dong:2020nwy , the transition amplitude can be written as
@@
where $\alpha$ represents an overall factor, $r_{i}$ and $\phi_{i}$ describe the relative production strength and phase angle for channel $i$, respectively, and $\beta$ is used to simulate the background.
@@
We primarily focus on the CMS data, utilizing the LHCb data as supplementary due to the higher precision of the CMS muon detector. The fit parameters are as follows, excluding the semi-fixed overall factors $\alpha$ and $\beta$ which control the lineshape and background magnitude. We first select a set of $(\alpha,\beta)$ values, fix them, perform the fit, and then choose the best fit result for presentation. For each channel $i$, there are two parameters: the phase angle $\phi_{i}$ and the relative production strength $r_{i}$.
@@
For the di-$J/\psi$ invaraint mass spectrum, considering only the $S$-wave effective potential in the scattering, there are two possible quantum numbers: $0^{++}$ and $2^{++}$. The $1^{++}$ configuration is excluded because, although $1\otimes 1\rightarrow 1$ is possible, it cannot form a state with $C=+$ parity under $S$-wave conditions. Tab. 1 presents the results of the two sets of fit parameters for the $0^{++}$ and $2^{++}$ configurations. The corresponding fit results, based on the parameters in Tab. 1, are shown in Fig. 3.

### v2

# Discovering an unquenched dynamics mechanism for charmonium scattering

@@ Abstract
In this work, we propose a new mechanism for charmonium scattering that utilizes the internal structure of charmonium. By capturing light flavor quarks and anti-quarks from the vacuum, charm and anti-charm quarks form virtual charmed mesons, which mediate an effective one-boson exchange process. This approach accurately reproduces the di-$J/\psi$ invariant mass spectrum observed by CMS and LHCb, demonstrating its validity. Our mechanism offers a comprehensive framework for understanding charmonium scattering and is applicable to the scattering problems involving all fully heavy hadrons, an area of increasing interest.
I Introduction
@@
Scattering problems, as a fundamental topic in physics, offering critical insights into the interactions and internal structures of matter. Using this approach, Yukawa predicted the pion meson Yukawa:1935xg , which mediates the nuclear force, establishing a cornerstone of modern nuclear physics. In this century, advances in experimental techniques have led to the discovery of numerous new hadronic states, presenting unique opportunities to explore hadron scattering. Noteworthy achievements include the identification of exotic molecular states such as $P_{c}$ states LHCb:2015yax ; LHCb:2019kea and $T_{cc}(3875)^{+}$  LHCb:2021vvq ; LHCb:2021auc . These discoveries not only mark the advent of ’Particle 2.0’ but also enhance our understanding of the nonperturbative aspects of the strong force Chen:2016qju ; Richard:2016eis ; Chen:2016spr ; Lebed:2016hpi ; Esposito:2016noz ; Guo:2017jvc ; Ali:2017jda ; Brambilla:2019esw ; Liu:2019zoy ; Chen:2022asf .
@@
Recently, the CMS Collaboration reported a high-precision measurement of the invariant mass spectrum of di-$J/\psi$ CMS:2023owd , confirming the previous observation of $X(6900)$ by the LHCb Collaboration LHCb:2020bwg and revealing additional enhancements in the di-$J/\psi$ invariant mass spectrum. Furthermore, ATLAS has joined this effort, alongside LHCb and CMS ATLAS:2023bft . This new finding surprised the community, prompting various explanations, including exotic hadronic states liu:2020eha ; Tiwari:2021tmz ; Lu:2020cns ; Faustov:2020qfm ; Zhang:2020xtb ; Li:2021ygk ; Liu:2021rtn ; Giron:2020wpx ; Karliner:2020dta ; Wang:2020ols ; Ke:2021iyh ; Zhu:2020xni ; Jin:2020jfc ; Yang:2021hrb ; Albuquerque:2020hio ; Albuquerque:2021erv ; Wu:2022qwd ; Asadi:2021ids ; Yang:2020wkh ; Chen:2020xwe ; Deng:2020iqw ; Gordillo:2020sgc ; Zhao:2020zjh ; Faustov:2021hjs ; Wang:2021kfv and coupled-channel dynamics generation Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . However, a substantial issue—charmonium scattering—has not been fundamentally resolved, despite some tentative efforts Wang:2020wrp ; Wang:2020tpt ; Wang:2022jmb ; Gong:2020bmg ; Dong:2020nwy ; Dong:2021lkh ; Liang:2021fzr ; Lu:2023aal ; Guo:2020pvt ; Zhuang:2021pci . The highly effective one-boson ($\pi$, $\sigma$, $\rho$, $\cdots$) exchange mechanism for describing hadron scattering fails in quantitatively characterizing charmonium scattering, as charmonium is a typical example of a fully heavy hadron. Thus, studying how charmonium undergoes scattering has become a particularly typical issue in current scattering problems and urgently requires a solution.
@@
Figure 1: A new approach enabling one-boson exchange for effective charmonium scattering. The cross symbol on the left indicates that direct light flavor one-boson exchange ($\pi,\rho,\sigma,$$\cdots$) cannot occur.
@@
In this work, we propose a new mechanism for charmonium scattering. As illustrated in Fig. 1, the charm and anti-charm quarks within charmonium can capture light flavor anti-quark and quark, respectively, created from the vacuum to form virtual charmed mesons. These virtual mesons provide a source that enables one-boson exchange to be effective for charmonium scattering. A direct application of this new mechanism successfully reproduces the di-$J/\psi$ invariant mass spectrum data from CMS and LHCb, demonstrating the mechanism’s validity. We should emphasize that this mechanism is not limited to charmonium scattering; it can also be applied to investigate the scattering problems involving all fully heavy hadrons, a topic that is gaining attention from the community, including lattice research groups Lyu:2021qsh ; Mathur:2022ovu .
II Application to reproduce the CMS and LHCb data
@@
The di-$J/\psi$ invariant mass spectrum observed by the CMS and LHCb Collaborations exhibit significant features around the dip near 6.8 GeV and the peak near 6.9 GeV, corresponding closely to the $J/\psi\psi(3686)$ and $J/\psi\psi(3770)$ thresholds, respectively. In contrast, the bumps near 6.6 GeV and 7.3 GeV do not align with any thresholds involving particles from the $\psi$ family. Therefore, to capture the main features, we consider couplings among three channels: $J/\psi J/\psi,J/\psi\psi(3686),J/\psi\psi(3770)$}. As for the di-$\eta_{c}$ and di-$\chi_{cJ}$ channels, although there may exist thresholds that close to the observed peaks, due to their much lower production cross sections Bodwin:1994jh ; ATLAS:2015zdw ; ATLAS:2014ala ; CMS:2012qwg ; LHCb:2013ofo ; He:2024lrb , we neglect their effects in this work.
@@
When one-boson exchange is introduced at the quark level, as depicted in Fig. 1, the scattering mechanism between two charmonia can proceed as shown in Fig. 2. Initially, the two charmonia, $\psi_{1}$ and $\psi_{2}$, couple into pairs of virtual $D$ mesons. Within each pair, one $D$ meson acts as a spectator, while the other engages in an interaction mediated by the exchange of a light hadron, $\mathcal{P}$. Subsequently, the interacting virtual $D$ mesons combine with the spectators to form the final charmonia, $\psi_{3}$ and $\psi_{4}$
@@
As shown in Fig. 2, we have neglected the contributions from charmed-strange meson loops. This is primarily due to the significantly lower creation rate of $s\bar{s}$ quark pairs compared to $u\bar{u}$ or $d\bar{d}$ quark pairs, with a ratio of approximately ${1}/{\sqrt{3}}$. Additionally, the higher threshold for producing charmed-strange meson pairs leads to further suppression on loop function. Therefore, to focus on the primary physics, we exclude contributions from charmed-strange meson loops.
Furthermore, Fig. 2 illustrates that the exchanged light hadron, $\mathcal{P}$, is limited to $\sigma$ and $\eta$. This restriction arises because the final interaction between two charmonia and a light hadron still must comply with the conservation laws of the strong interaction. Given that the mass of the exchanged particle should be as small as possible, only $\sigma$ and $\eta$ are viable candidates, they can contribute the intermediate and long range interactions, whereas, the direct charmed mesons exchanges provide a very short range interactions due to their large masses Yukawa:1935xg , although it seems to be more efficient interactions to study the charmonia interactions. Our result demonstrates that within our model the sigma-exchange dominates over the eta-exchange, as expected [^1] With the fit parameters listed in Table 1, we calculate the ratio of the $|\mathcal{M}^{J}_{J/\psi J/\psi}(\sqrt{s})|^{2}$, defined in Eq.(II), that purely caused by $\sigma$ and $\eta$ exchanges. The numerical showed that for both $0^{++}$ and $2^{++}$ cases, when $\sqrt{s}$ varies from 6.3 to 9.0 GeV, this ratio is always much larger than one..
We perform the entire calculation via the effective Lagrangian approach, with relevant Lagrangians being Cheng:1992xi ; Yan:1992gz ; Wise:1992hn ; Burdman:1992gh ; Casalbuoni:1996pg
@@
Here, $v$ is the velocity of heavy quark. $g=0.59$, $f_{\pi}=132$ MeV, and $g_{s}=0.76$ are the fixed coupling constants Ding:2021igr ; Falk:1992cx ; Isola:2003fh ; Liu:2009qhy ; Chen:2019asm . $g_{D^{\ast}D^{\ast}\psi}$, $g_{DD\psi}$, and $g_{D^{\ast}D\psi}$ can be related to one single parameter $f_{\psi}$ as $\frac{g_{D^{\ast}D^{\ast}\psi}}{m_{D^{\ast}}}=\frac{g_{DD\psi}}{m_{D}}=g_{D^{\ast}D\psi}=\frac{m_{\psi}}{m_{D}f_{\psi}}$ due to heavy quark effective theory.
@@
$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
@@
$\times\int\frac{d^{4}q_{1}}{(2\pi)^{4}}\frac{\mathcal{V}_{\mu\alpha}}{\prod_{i}(p_{i}^{2}-m_{i}^{2}+im_{i}\Gamma_{i})}\mathcal{F}_{q_{1}}(q_{1})$
$\times\int\frac{d^{4}q_{2}}{(2\pi)^{4}}\frac{\mathcal{V}_{\nu\beta}}{\prod_{j}(p_{j}^{2}-m_{j}^{2}+im_{j}\Gamma_{j})}\mathcal{F}_{q_{2}}(q_{2}),$
@@
where $\mathcal{V}_{\gamma\delta}$ denotes the Lorentz structure of the loop integral, $\mathcal{F}_{q}(q)$ is the form factor for the exchanged light hadron $\mathcal{P}$ to account for the off-shell effect, and $\mathcal{F}_{q_{i}}(q_{i})$ is a form factor used to avoid ultraviolet divergence in the loop integral. We use the widely adopted monopole form for $\mathcal{F}_{q}(q)$ as $\mathcal{F}_{q}(q)=\frac{\Lambda_{q}^{2}-m_{q}^{2}}{\Lambda_{q}^{2}-q^{2}}$, where $\Lambda_{q}$ is parametrized as $\Lambda_{q}=m_{q}+\alpha_{q}\Lambda_{\mathrm{QCD}}$, with $\Lambda_{\mathrm{QCD}}=220$ MeV and $\alpha_{q}$ is an adjustable parameter. For $\mathcal{F}_{q_{i}}(q_{i})$, we adopt $\mathcal{F}_{q_{i}}(q_{i})=\exp(-\vec{q}_{i}^{~2}/\Lambda_{i}^{2})$.
For both Eq. (6) and Eq. (7), we multiply them with an overall factor of 16. It is because each $D$ meson loop in Fig. 2 can involve not only $D^{(\ast)}D^{(\ast)}\bar{D}^{(\ast)}$, but also $\bar{D}^{(\ast)}\bar{D}^{(\ast)}D^{(\ast)}$ configurations. Furthermore, for both of these cases, there have charged and neutral situations. Thus, each loop has a total of 4 configurations, resulting in 16 configurations for each diagram in Fig. 2. We want to emphasize here that for $J/\psi J/\psi$ interactions in this model, although it seems the diagram may be suppressed by the masses of the exchanged mesons, like $1/(m_{J/\psi}^{2}-4m_{D}^{2})\times 1/m_{\mathcal{P}}^{2}$, this suppression can be overcame by the summations on the amplitudes with all the possible intermediate $D^{(\ast)}$ and $\bar{D}^{(\ast)}$ combinations in addition with a large but not unreasonable cutoff in form factor. In particular, when the discussed charmonia has a larger mass, the two $D$-mesons connected with it will be more close to their mass shells, which can reduce the suppression factor significantly.
@@
In this work, we consider only the $S$-wave behavior in the di-$J/\psi$ spectrum. Therefore, we first perform a partial wave expansion on the amplitude to get the interaction kernel, which can be calculated as
$\mathcal{K}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})=2\pi\int d\cos\theta d_{\lambda\lambda^{\prime}}^{J}(\theta)\mathcal{M}_{\lambda,\lambda^{\prime}}(\sqrt{s},\vec{k},\vec{k}^{\prime})$ (8)
@@
with $k^{(\prime)}=|\vec{k}^{~(\prime)}|$ being the module of the three-momenta of initial and final particles in the center-of-mass system, $\lambda$ and $\lambda^{\prime}$ is the helicity indices of the initial and final states, respectively, and $d_{\lambda,\lambda^{\prime}}^{J}(\theta)$ is the Wigner-$d$ matrix with $\theta$ being the relative angle between $\vec{k}$ and $\vec{k}^{\prime}$.
@@
Next, we substitute the partial wave interaction kernel into the coupled-channel Lippmann-Schwinger equation
$T^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})$ $=$ $\mathcal{K}^{J}_{\lambda,\lambda^{\prime}}(\sqrt{s},k,k^{\prime})+\sum_{\lambda^{\prime\prime}}\int\frac{k^{\prime\prime 2}dk^{\prime\prime}}{(2\pi)^{3}}$ (9)
$\times\mathcal{K}^{J}_{\lambda,\lambda^{\prime\prime}}(\sqrt{s},k,k^{\prime\prime})G_{0}(\sqrt{s},k^{\prime\prime})$
$\times T^{J}_{\lambda^{\prime\prime},\lambda^{\prime}}(\sqrt{s},k^{\prime\prime},k^{\prime}),$
@@
then, with the extracted on-shell coupled-channel $T$-matrices $T^{J}(\sqrt{s})$, the transition amplitude can be written as Gong:2020bmg ; Dong:2020nwy
@@
where $G_{0}$ is the free two particle Green’s function, $G_{0i}$ means the free Green’s function of $i$-th channel, $\alpha$ represents an overall factor, $r_{i}$ and $\phi_{i}$ describe the relative production strength and phase angle for channel $i$, respectively, and $\beta$ is used to simulate the background. Although the introduction of $r_{i}$ and $\phi_{i}$ will violate the unitary, however, unitarity is strictly maintained only within the scattering sector—calculated via the Lippmann–Schwinger equation for the coupled channels $\{J/\psi J/\psi,J/\psi\psi(3686),J/\psi\psi(3770)\}$. When this unitary scattering amplitude is combined with the parameterized, non-calculated production amplitude to form the full observable amplitude, overall unitarity is broken, as we do not perform a fully coupled-channel calculation including the $pp$ state.
@@
We primarily focus on the CMS data, utilizing the LHCb data as supplementary due to the higher precision of the CMS muon detector. The fit parameters are as follows, excluding the semi-fixed overall factors $\alpha$ and $\beta$ which control the lineshape and magnitude of background, for each channel $i$, there are two parameters: the phase angle $\phi_{i}$ and the relative production strength $r_{i}$.
@@
For the di-$J/\psi$ invaraint mass spectrum, considering only the $S$-wave effective potential in the scattering, there are two possible quantum numbers: $0^{++}$ and $2^{++}$. The $1^{++}$ configuration is excluded because, although $1\otimes 1\rightarrow 1$ is possible, it cannot form a state with $C=+$ parity under $S$-wave conditions. Table 1 presents the results of the two sets of fit parameters for the $0^{++}$ and $2^{++}$ configurations. The corresponding fit results extracted pole positions, based on the parameters in Tab. 1, are shown in Fig. 3 and Tab. 2, respectively.
@@
Table 2: The pole positions $m-\frac{\Gamma}{2}i$ extracted from our fit parameters given in Table 1 for the $0^{++}$ and $2^{++}$ cases. All values are given in units of MeV. Here, the uncertainties in the pole positions come from the variations of the cutoffs within 50 MeV.
$0^{++}$ case $2^{++}$ case
$(6204\pm 24)-(23\pm 23)i$ $(6198\pm 7)-(7\pm 7)i$
$(6789\pm 17)-(51\pm 51)i$ $(6783\pm 9)-(8\pm 8)i$
$(6871\pm 11)-(25\pm 25)i$ $(6871\pm 10)-(38\pm 38)i$